a spherical capacitor has an inner sphere radius 10cm and an outer sphere of radius 14cm. the outer sphere is earthed and the onner sphere is given a charge of 2.5"uC" . the space between the concentric spheres is filled with a liquid of dielectric constant 32. determine the capacitamce of the capacitor.


1
Expert's answer
2021-04-28T13:04:48-0400
"C=\frac{4\pi \epsilon \epsilon_0 rR}{R-r}\\
C=\frac{4\pi (32)(8.85\cdot10^{-12})(0.14)(0.1)}{0.14-0.1}=1.25\cdot10^{-9}\ F"